Given ; Two dices are rolled


To Find ; probability of getting similar faces
Answer #1

Solution ; probability of an event be find out using the formula =


number of favourable outcomes/ total number of possible outcomes


according to question we have to take case of similar ones


possible cases = 6


Total cases =36


probability =6/36 =1/6


Hence the probability is 1/6


answered by: anonymous